Find the pressure difference on an airplane wing if air flows over the upper surface with a speed of 115 m/s , and along the bottom surface with a speed of 103 m/s . Answer is in kPa.

If the area of the wing is 33 m2 , what is the net upward force exerted on the wing?

Express your answer using two significant figures.

Answer in kN
Answer #1

Here ,

density of air , p = 1.225 Kg/m^3

let the pressure difference is P

Using Bernoulli's equation

P = 0.5 * p * (v1^2 - v2^2)

P = 0.5 * 1.225 * (115^2 - 103^2)

P = 1602.3 Pa

P = 1.602 kPa

the pressre difference on an airplane wing is 1.602 kPa

net upward force = P * Area

net upward force = 1602 * 33

net upward force = 52875.9 N = 52.88 kN

the new upward force on the wings is 52.88 kN